# Invertible Calabi-Yau Orbifolds over Finite Fields

### v1

@@ 1. Introduction
In this paper we continue the line of inquiry started in [AP] and further investigate the arithmetic information stored in the $p$-adic deformed Borisov complex. Both the Borisov complex and its deformation introduced in [AP] are constructed out of polynomials in two sets of variables, corresponding to the variables of the superpotentials on the two mirror models. While the Frobenius operator used in [AP] acts on all the variables, in this paper we restrict it to only one set of variables. Additionally, we replace the standard Frobenius operator with its more arithmetically natural inverse (see e.g. [M]). Finally, following earlier work on the (étale) cohomology of orbifolds [Behrend93, Rose07], we twist the Frobenius operator by a suitable power of $p$ depending on the notion of age. Our first observation is that if the determinant of the matrix defining the superpotential divides $p-1$, then our modified Frobenius operator is diagonalized by the standard basis of the total cohomology of the deformed Borisov bicomplex. Furthermore, we explicitly express the eigenvalues of the cohomological Frobenius in terms of $p$-adic gamma functions.
Drawing Motivation from the LG/CY correspondence and from methods for calculating hypersurface point counts over finite fields via traces of the Frobenius endomorphism acting on the Monsky-Washnitzer cohomology, we form a conjecture: the number of points of any crepant resolution of a Berglund-Hübsch orbifold can be computed by the supertrace of our modified Frobenius operator acting on the cohomology of the deformed $p$-adic Borisov complex associated to the orbifold. As evidence for our conjecture, we first point out that it correctly counts points modulo $p$. We then prove the conjecture for all Berglund-Hübsch elliptic curves (including those expressed in terms of chains, loops, and their combinations), and for all weighted diagonal K3 surfaces. In the latter case, our orbifold calculations recover the results obtained by Goto [Goto] through a careful analysis of the corresponding minimal resolutions of singularities.
@@ 2. Preliminaries
is invertible; i.e. quasi-homogeneous and such that $(\partial_{x_{1}}W_{A},\ldots\partial_{x_{n}}W_{A})$ is a regular sequence in $\mathbb{F}_{p}[x]=\mathbb{F}_{p}[x_{1},\ldots,x_{n}]$. Here $e_{1},e_{2},\ldots,e_{n}$ are the elements of the standard basis of $\mathbb{Z}^{n}$. We also assume that $\det(A)$ divides $p-1$.
@@
$A$ is Berglund-Hübsch over $\mathbb{F}_{p}$ if and only if $A^{T}$ is Berglund-Hübsch over $\mathbb{F}_{p}$.
@@
Let $A$ be Berglund-Hübsch matrix $A$ over $\mathbb{F}_{p}$. Let ~$\widetilde{\mathcal{B}}_{A}$ be the subspace of overconvergent series in the even variables $\{x_{i},y_{i}\}_{i=1}^{n}$ and in the odd variables $\{\theta_{i}\}_{i=1}^{n}$ spanned by all monomials $x^{\gamma}y^{\lambda}\theta^{I}$ with the property that $\lambda A^{-T}\geq 0$. Let $\mathcal{B}_{A}$ be the quotient of ~$\widetilde{\mathcal{B}}_{A}$ by the ideal generated by the monomials $x^{\gamma}y^{\lambda}\theta^{I}$ such that $\gamma A^{-1}\lambda^{T}>0$. Let $Q_{A}^{\vee}$ be the grading operator acting on $\mathcal{B}_{A}$ with respect to which each monomial $x^{\gamma}y^{\lambda}\theta^{I}$ is an eigenvector with eigenvalue $q^{\vee}(\lambda,I)$ equal to the number of indices $i\in\{1,\ldots,n\}$ such that $(\lambda A^{-T})_{i}>(\lambda A^{-T})_{i}I_{i}$. Similarly, the dual grading operator $Q_{A}$ has the property that each monomial $x^{\gamma}y^{\lambda}\theta^{I}$ is an eigenvector with eigenvalue
@@
    $(\mathcal{B}_{A},d_{A},d_{A}^{\vee})$ decomposes into the direct sum of sub-bicomplexes $(\mathcal{B}_{A}(\gamma,\lambda),d_{A},d_{A}^{\vee})$ labeled by classes $[(\gamma,\lambda)]\in\mathbb{Z}^{2n}/(\mathbb{Z}^{2n}(A\oplus A^{T}))\cong G_{A^{T}}\times G_{A}$.
@@
Let $A$ be a Berglund-Hübsch matrix over $\mathbb{F}_{p}$. The transpose of a subgroup $G$ of $G_{A}$ is the subgroup of $G^{T}$ consisting of all $\gamma\in G_{A^{T}}$ such that $\gamma A^{-1}\lambda^{T}\in\mathbb{Z}$ for all $\lambda\in G$.
@@
Let $A$ be an $n\times n$ Berglund-Hübsch matrix over $\mathbb{F}_{p}$. Assume the representative in the equivalence class of $\lambda\in G_{A}$ modulo $\mathbb{Z}^{n}A^{T}$ is chosen so that $10\leq(\lambda A^{-T})_{i}<$ 1 for every $i\in\{1,\ldots,n\}$. The age of $\lambda$ is ${\rm age}(\lambda)=\sum_{i=1}^{n}(\lambda A^{-T})_{i}$. Similarly, the dual age of $\gamma\in G_{A^{T}}$ is defined to be ${\rm age}^{\vee}(\gamma)=\sum_{i=1}^{n}(\gamma A^{-1})_{i}$. We also define $\dim(\lambda)$ to be the number of indices $i$ for which $(\lambda A^{-T})_{i}\in\mathbb{Z}$.
@@ 3. Frobenius Maps
where $s\in\mathbb{N}^{n}$ and $[-\gamma A^{-1}]\in\mathbb{N}^{n}$ is the vector satisfying $0\leq[-\gamma A^{-1}]_{i}<p$ and $[-\gamma A^{-1}]A\equiv-\gamma\mod{p}$. This lets us rewrite the sum in terms of $s$. In particular, define $\alpha$ so that
@@
By [A], the $p$-adic gamma function satisfies
@@
${\rm ST}_{p}(A,G)\equiv 1+(-1)^{n-1}\Gamma_{p}(JA^{-1})\mod p$ (3.21)
. Using the reflection formula for the $p$-adic gamma function in the form
@@
valid whenever $d$ is an integer dividing $p-1$ and $a\in\{1,\ldots,d-1\}$, and the fact that $\Gamma_{p}(x)$ is congruent mod $p$ to $\Gamma_{p}(y)$ whenever $x$ and $y$ are [Koblitz80], we obtain
11 11 $=(-1)^{p-(p-1)(JA^{-1})_{i}}\Gamma_{p}((JA^{-1})_{i})\Gamma_{p}(1-(JA^{-1})_{i})$
@@
Since $(p-1)JA^{-1}$ is an integer vector, we see from the product representation [Koblitz80] of $\Gamma_{p}(p-(p-1)(1-(JA^{-1})_{i}))$ that
@@
#$\#X_{A}(\mathbb{F}_{p})$ $\equiv 1+\sum_{x\in\mathbb{F}_{p}^{n}}(W_{A}(x))^{p-1}\,\mod p$
@@ 4. An Arithmetic Conjecture
Let $A$ be an $n\times n$ Berglund-Hübsch matrix over $\mathbb{F}_{p}$, let $n\geq 3$, and let $G$ be a subgroup of $G_{A}$ such that $J\in G\cap G^{T}$. Assume that $JA^{-T}J^{T}=1$ and let $X_{A}$ be the hypersurface in the weighted projective space over $\mathbb{F}_{p}$ with weights $w_{i}=m(JA^{-T})_{i}$, where $m$ is the smallest integer such that each $w_{i}$ is an integer, cut out by $W_{A}(x)=0$. If $Y_{A,G}$ is the orbifold $[X_{A}/\langle J\rangle]$, then the number of points of any crepant resolution $Z_{A,G}(\mathbb{F}_{p})$ of $Y_{A,G}(\mathbb{F}_{p})$ is equal to ${\rm ST}_{p}(A,G)$.
@@
Wan’s congruence mirror conjecture [Wan05] implies that if $(X,X^{\prime})$ is a strong mirror pair of Calabi-Yau manifolds defined over $\mathbb{F}_{p}$, then #$\#X(\mathbb{F}_{p})$ is congruent to #$\#X^{\prime}(\mathbb{F}_{p})$ modulo $p$. As pointed out in [Wan05], the notion of a strong mirror pair lacks a rigorous general definition. In [MagyarWhitcher18], a strong mirror pair is defined as a mirror pair for which the congruence mirror conjecture holds. In any case, the congruence mirror conjecture has been proven [W] for Dwork pencils and their mirrors. It is natural to ask, as done in [MagyarWhitcher18, DKSSVW18, Whitcher21, SalernoWhitcher22], whether the Berglund-Hübsch construction [BH] is compatible with the congruence mirror conjecture – more precisely, whether #$\#Z_{A,G}(\mathbb{F}_{p})$ is congruent to #$\#Z_{A^{T},G^{T}}(\mathbb{F}_{p})$ modulo $p$.
Under the assumptions of Corollary 3.7, the congruence class modulo $p$ of ${\rm ST}_{p}(A,G)$ is independent of $G$. This means that Conjecture 4.1 implies the congruence mirror conjecture when $A$ is diagonal. On the other hand, if $A$ is not diagonal, there is no reason to expect ${\rm ST}_{p}(A,G)$ and ${\rm ST}_{p}(A^{T},G^{T})$ to be congruent modulo $p$. Indeed, this is false for $W_{A}(x)=x_{1}^{2}x_{2}+x_{2}^{3}+x_{3}^{3}$ and $p=19$: using (3.21), ${\rm ST}_{19}(A,\langle J\rangle)\equiv 9\mod 19$ and ${\rm ST}_{19}(A^{T},\langle J\rangle)\equiv 8\mod 19$. In other words, Corollary 3.7 implies that the congruence mirror conjecture does not apply to the pair of Berglund-Hübsch mirror elliptic curves $Z_{A,\langle J\rangle}=X_{A}$ and $Z_{A^{T},\langle J\rangle}=X_{A^{T}}$.
@@
where $t^{\prime}$ is chosen to satisfy $t^{\prime}\equiv t\text{ mod }p-1$ and $10<t^{\prime}\leq p-$ 1. Adapting the calculations of [Weil49] to the $p$-adic setting by employing the Techmuüller and Dwork characters introduced above, we have that the number of points of $Z_{A,G}(\mathbb{F}_{p})$ is equal to
@@
valid whenever $d$ divides $p-1$, and identifying indices by setting $\beta_{i}=A_{ii}-\gamma_{i}$, we obtain the equality of (4.2) and (4.4). ∎
@@
Table 1. Singular weighted diagonal K3 hypersurfaces. Multiplicity refers to the number of singular points of that type on $X_{A}$ The last column lists the number of $-2$ curves appearing in the minimal resolution of the corresponding singularity
@@
Weighted diagonal K3 surfaces over finite fields are classified in [Goto]. Two of them, the Fermat quartic in $\mathbb{P}^{3}$ and the hypersurface $x_{1}^{2}+x_{2}^{6}+x_{3}^{6}+x_{4}^{6}=0$ in $\mathbb{P}(3,1,1,1)$ are smooth and Theorem 4.3 applies to them. The remaining $12$ cases with corresponding lists of singularities and the number of $-2$ curves added in the corresponding minimal resolution are listed in Table 1. In general, as observed in [Goto], some of the singularities in Table 1 are only defined over suitable field extensions. However, because in our case the coefficients in front of the monomials of $W_{A}(x)$ are always equal to $1$, these extensions are generated by elements $\gamma$ such that $01+\gamma^{m}=$ 0 with $m\in\{1,2\}$ depending on the singularity. On the other hand, inspection of Table 1 shows that $\det(A)$, and thus $p-1$, is always divisible by $4$. Hence our assumptions guarantee that $\mathbb{F}_{p}$ has primitive fourth roots of unity and thus $\mathbb{F}_{p}(\gamma)=\mathbb{F}_{p}$. Therefore, all the singularities in Table 1 actually occur, just as they do over $\mathbb{C}$, when $A$ satisfies the assumptions of Conjecture 4.1. Combining the calculations of [Weil49] with the (4.5) as in the proof of Theorem 4.3, we obtain
#$\#X_{A}(\mathbb{F}_{p})=1+p+p^{2}+\sum_{\gamma}(-p)^{{\rm age}^{\vee}(\gamma)-1}\Gamma_{p}(\gamma A^{-1})\,,$ (4.9)
@@
#$\#Z_{A,\langle J\rangle}(\mathbb{F}_{p})=1+(\nu+1)p+p^{2}+\sum_{\gamma}(-p)^{{\rm age}^{\vee}(\gamma)-1}\Gamma_{p}(\gamma A^{-1})\,,$ (4.10)
where $\nu$ is the total number of $-2$ curves appearing in $Z_{A,\langle J\rangle}$ viewed as the minimal resolution of $X_{A}$. Since the sum over $\gamma$ in (4.10) is equal to the contribution from terms labeled by $\lambda=0$ terms to (3.18), it remains to check that the contribution from the $\lambda\neq 0$ terms in (3.18) is exactly $1+(\nu+1)p+p^{2}$. It is easy to see that $y_{1}y_{2}y_{3}y_{4}$ contributes to 1, while $y_{1}^{A_{11}-1}$ where $\nu$ is the total number of $-2$ curves appearing in $Z_{A,\langle J\rangle}$ viewed as the minimal resolution of $X_{A}$. Since the sum over $\gamma$ in (4.10) is equal to the contribution from terms labeled by $\lambda=0$ terms to (3.18), it suffices to check that the contribution from the $\lambda\neq 0$ terms in (3.18) is exactly $1+(\nu+1)p+p^{2}$. It is easy to see that $y_{1}y_{2}y_{3}y_{4}$ contributes $1$ while $y_{1}^{A_{11}-1}y_{2}^{A_{22}-1}y_{3}^{A_{33}-1}y_{4}^{A_{44}-1}$ contributes $p^{2}$. By Remark 2.22, we know that the remaining $\lambda\neq 0$ terms all correspond to cohomology classes in $H^{1,1}(Z_{A,G}(\mathbb{C}))$. In particular, there must be precisely $\nu$ of them and they must be of one of two kinds: monomials of the form $y^{\lambda}$ with ${\rm age}(\lambda)=1$ and monomials of the form $x_{i}^{\gamma_{i}}x_{j}^{\gamma_{j}}y_{r}^{\lambda_{r}}y_{s}^{\lambda_{s}}\theta_{i}\theta_{j}$ with ${\rm age}(\lambda)=1={\rm age}^{\vee}(\gamma)$. Each of these $y^{\lambda}$ monomials contributes $p$ to (3.18). The constraint ${\rm age}^{\vee}(\gamma)=1$ forces each monomial $x_{i}^{\gamma_{i}}x_{j}^{\gamma_{j}}y_{r}^{\lambda_{r}}y_{s}^{\lambda_{s}}\theta_{i}\theta_{j}$ to be such that $\frac{\gamma_{i}}{A_{ii}}+\frac{\gamma_{j}}{A_{jj}}=1$, making the corresponding contribution to (3.18) equal to
$-p\Gamma_{p}\left(\frac{\gamma_{i}}{A_{ii}}\right)\Gamma_{p}\left(\frac{\gamma_{j}}{A_{jj}}\right)=-p(-1)^{p-(p-1)\frac{\gamma_{i}}{A_{ii}}}\,,$ (4.11)
where the reflection formula (3.22) was used to derive (4.11). Inspection of Table 1 shows that each $W_{A}(x)$ has at least two terms with even exponents. Taking again into account that $\det(A)$ divides $p-1$, we conclude that $\frac{p-1}{A_{ii}}$ is necessarily even and thus $p-(p-1)\frac{\gamma_{i}}{A_{ii}}$ is necessarily odd. Hence, monomials of the form $x_{i}^{\gamma_{i}}x_{j}^{\gamma_{j}}y_{r}^{\lambda_{r}}y_{s}^{\lambda_{s}}\theta_{i}\theta_{j}$ contribute by $p$ to (3.18), concluding the proof.
@@
$\left(0,0\right)$ $y_{1}^{1}y_{2}^{1}y_{3}^{1}y_{4}^{1}$ $1$
$\left(2,0\right)$ $x_{1}^{1}x_{2}^{1}x_{3}^{1}x_{4}^{1}\theta_{1}\theta_{2}\theta_{3}\theta_{4}$ $-\,\Gamma_{p}\left(\frac{1}{2}\right)\Gamma_{p}\left(\frac{1}{3}\right)\Gamma_{p}\left(\frac{1}{10}\right)\Gamma_{p}\left(\frac{1}{15}\right)$

### v2

@@ 1. Introduction
In this paper we continue the line of inquiry started in [AP] and further investigate the arithmetic information stored in the $p$-adic deformed Borisov complex. Both the Borisov complex and its deformation introduced in [AP] are constructed out of polynomials in two sets of variables, corresponding to the variables of the superpotentials on the two mirror models. While the Frobenius operator used in [AP] acts on all the variables, in this paper we restrict it to only one set of variables. Additionally, we replace the standard Frobenius operator with its more arithmetically natural inverse (see e.g. [M]). Finally, following earlier work on the (étale) cohomology of orbifolds [Behrend93, Rose07], we twist the Frobenius operator by a suitable power of $p$ depending on the notion of age. Our first observation is that if the determinant of the matrix defining the superpotential divides p1p-1, then our modified Frobenius operator is diagonalized by the standard basis of the total cohomology of the deformed Borisov bicomplex. Furthermore, we explicitly express the eigenvalues of the cohomological Frobenius in terms of $p$-adic gamma functions.
Drawing motivation from the LG/CY correspondence [ChiodoRuan11] and from methods for calculating hypersurface point counts over finite fields via traces of the Frobenius endomorphism acting on the Monsky-Washnitzer cohomology, we form a conjecture: the number of points of any crepant resolution of a Berglund-Hübsch orbifold can be computed by the supertrace of our modified Frobenius operator acting on the cohomology of the deformed $p$-adic Borisov complex associated to the orbifold. As evidence for our conjecture, we first point out that it correctly counts points modulo $p$. We then prove the conjecture for all Berglund-Hübsch elliptic curves (including those expressed in terms of chains, loops, and their combinations), and for all weighted diagonal K3 surfaces. In the latter case, our orbifold calculations recover the results obtained by Goto [Goto] through a careful analysis of the corresponding minimal resolutions of singularities.
@@ 2. Preliminaries
is invertible; i.e. quasi-homogeneous and such that $(\partial_{x_{1}}W_{A},\ldots\partial_{x_{n}}W_{A})$ is a regular sequence in $\mathbb{F}_{p}[x]=\mathbb{F}_{p}[x_{1},\ldots,x_{n}]$. Here $e_{1},e_{2},\ldots,e_{n}$ are the elements of the standard basis of $\mathbb{Z}^{n}$ (throughout the paper, we assume our vectors in $\mathbb{Z}^{n}$ to be row vectors). We also assume that $\det(A)$ divides p1p-1.
@@
AA is Berglund-Hübsch over $\mathbb{F}_{p}$ if and only if $A^{T}$ is Berglund-Hübsch over $\mathbb{F}_{p}$.
@@
Let $A$ be Berglund-Hübsch matrix $A$ over $\mathbb{F}_{p}$ and let $A^{-T}$ be its inverse transpose. Let $\widetilde{\mathcal{B}}_{A}$ be the subspace of overconvergent series in the even variables $\{x_{i},y_{i}\}_{i=1}^{n}$ and in the odd variables $\{\theta_{i}\}_{i=1}^{n}$ spanned by all monomials $x^{\gamma}y^{\lambda}\theta^{I}$ with the property that $\lambda A^{-T}\geq 0$. Let $\mathcal{B}_{A}$ be the quotient of $\widetilde{\mathcal{B}}_{A}$ by the ideal generated by the monomials $x^{\gamma}y^{\lambda}\theta^{I}$ such that $\gamma A^{-1}\lambda^{T}>0$. Let $Q_{A}^{\vee}$ be the grading operator acting on $\mathcal{B}_{A}$ with respect to which each monomial $x^{\gamma}y^{\lambda}\theta^{I}$ is an eigenvector with eigenvalue $q^{\vee}(\lambda,I)$ equal to the number of indices $i\in\{1,\ldots,n\}$ such that $(\lambda A^{-T})_{i}>(\lambda A^{-T})_{i}I_{i}$. Similarly, the dual grading operator $Q_{A}$ has the property that each monomial $x^{\gamma}y^{\lambda}\theta^{I}$ is an eigenvector with eigenvalue
@@
    $(\mathcal{B}_{A},d_{A},d_{A}^{\vee})$ decomposes into the direct sum of sub-bicomplexes $(\mathcal{B}_{A}(\gamma,\lambda),d_{A},d_{A}^{\vee})$ labeled by classes $(\gamma,\lambda)]\in\mathbb{Z}^{2n}/(\mathbb{Z}^{2n}(A\oplus A^{T}))\cong G_{A^{T}}\times G_{A}$.
@@
Let $A$ be a Berglund-Hübsch matrix over $\mathbb{F}_{p}$ and let $G$ be a subgroup of $G_{A}$. The transpose of $G$ is the subgroup $G^{T}$ of $G_{A^{T}}$ defined by requiring that an element $\gamma\in G_{A^{T}}$ belongs to $G^{T}$ if and only if $\gamma A^{-1}\lambda^{T}\in\mathbb{Z}$ for all $\lambda\in G$.
@@
Let $A$ be an $n\times n$ Berglund-Hübsch matrix over $\mathbb{F}_{p}$. Assume the representative in the equivalence class of $\lambda\in G_{A}$ modulo $\mathbb{Z}^{n}A^{T}$ is chosen so that $0\leq(\lambda A^{-T})_{i}<1$ for every $i\in\{1,\ldots,n\}$. The age of $\lambda$ is ${\rm age}(\lambda)=\sum_{i=1}^{n}(\lambda A^{-T})_{i}$. Similarly, the dual age of $\gamma\in G_{A^{T}}$ is defined to be ${\rm age}^{\vee}(\gamma)=\sum_{i=1}^{n}(\gamma A^{-1})_{i}$. We also define $\dim(\lambda)$ to be the number of indices $i$ for which $(\lambda A^{-T})_{i}\in\mathbb{Z}$.
@@ 3. Frobenius Maps
Let $\Gamma_{p}$ denote the $p$-adic gamma function. We refer the reader to [Koblitz84] and [Cohen07] for pedagogical accounts. In what follows, we find it helpful to employ the following notational convention. Given a vector $\alpha=(\alpha_{1},\ldots,\alpha_{n})\in\mathbb{Z}_{p}^{n}$, we denote by $\Gamma_{p}(\alpha)$ the product $\Gamma_{p}(\alpha_{1})\cdots\Gamma_{p}(\alpha_{n})$.
@@
where $s\in\mathbb{N}^{n}$ and $\gamma A^{-1}]\in\mathbb{N}^{n}$ is the vector satisfying $0\leq[-\gamma A^{-1}]_{i}<p$ and $\gamma A^{-1}]A\equiv-\gamma\mod{p}$. This lets us rewrite the sum in terms of $s$. In particular, define $\alpha$ so that
@@
By Proposition 11.6.15 in [Cohen07], the $p$-adic gamma function satisfies
@@
${\rm ST}_{p}(A,G)\equiv 1+(-1)^{n-1}\Gamma_{p}(JA^{-1})\mod p\,.$ (3.21)
Using the reflection formula for the $p$-adic gamma function in the form
@@
valid whenever $d$ is an integer dividing p1p-1 and $a\in\{1,\ldots,d-1\}$, and the fact that $\Gamma_{p}(x)$ is congruent mod $p$ to $\Gamma_{p}(y)$ whenever $x$ and $y$ are (See e.g. Sec. 2 Ch. 4 of [Koblitz80]), we obtain
$1$ $=(-1)^{p-(p-1)(JA^{-1})_{i}}\Gamma_{p}((JA^{-1})_{i})\Gamma_{p}(1-(JA^{-1})_{i})$
@@
Since $(p-1)JA^{-1}$ is an integer vector, we see from the product representation (ibid. ) of $\Gamma_{p}(p-(p-1)(1-(JA^{-1})_{i}))$ that
@@
$\#X_{A}(\mathbb{F}_{p})$ $\equiv 1+\sum_{x\in\mathbb{F}_{p}^{n}}(W_{A}(x))^{p-1}\,\mod p$
@@ 4. An Arithmetic Conjecture
Let $A$ be an $n\times n$ Berglund-Hübsch matrix over $\mathbb{F}_{p}$, let $n\geq 3$, and let $G$ be a subgroup of $G_{A}$ such that $J\in G\cap G^{T}$. Assume that $JA^{-T}J^{T}=1$ and let $X_{A}$ be the hypersurface in the weighted projective space over $\mathbb{F}_{p}$ with weights $w_{i}=m(JA^{-T})_{i}$, where $m$ is the smallest integer such that each $w_{i}$ is an integer, cut out by $W_{A}(x)=0$. If $Y_{A,G}$ is the orbifold $X_{A}/\langle J\rangle]$, then the number of points of any crepant resolution $Z_{A,G}(\mathbb{F}_{p})$ of $Y_{A,G}(\mathbb{F}_{p})$ is equal to ${\rm ST}_{p}(A,G)$.
@@
Wan’s congruence mirror conjecture [Wan05] implies that if $(X,X^{\prime})$ is a strong mirror pair of Calabi-Yau manifolds defined over $\mathbb{F}_{p}$, then $X(\mathbb{F}_{p})$ is congruent to $X^{\prime}(\mathbb{F}_{p})$ modulo $p$. As pointed out in [Wan05, WanFu06], the notion of a strong mirror pair lacks a rigorous general definition. In [MagyarWhitcher18], a strong mirror pair is defined as a mirror pair for which the congruence mirror conjecture holds. In any case, the congruence mirror conjecture has been proven [W] for Dwork pencils and their mirrors. It is natural to ask, as done in [MagyarWhitcher18, DKSSVW18, Whitcher21, SalernoWhitcher22], whether the Berglund-Hübsch construction [BH] is compatible with the congruence mirror conjecture – more precisely, whether $Z_{A,G}(\mathbb{F}_{p})$ is congruent to $Z_{A^{T},G^{T}}(\mathbb{F}_{p})$ modulo $p$.
Under the assumptions of Corollary 3.7, the congruence class modulo $p$ of ${\rm ST}_{p}(A,G)$ is independent of $G$. This means that Conjecture 4.1 implies the congruence mirror conjecture when $A$ is symmetric. On the other hand, if $A$ is not symmetric, there is no reason to expect ${\rm ST}_{p}(A,G)$ and ${\rm ST}_{p}(A^{T},G^{T})$ to be congruent modulo $p$. Indeed, this is false for $W_{A}(x)=x_{1}^{2}x_{2}+x_{2}^{3}+x_{3}^{3}$ and $p=19$: using (3.21), ${\rm ST}_{19}(A,\langle J\rangle)\equiv 9\mod 19$ and ${\rm ST}_{19}(A^{T},\langle J\rangle)\equiv 8\mod 19$. In other words, Corollary 3.7 implies that the congruence mirror conjecture does not apply to the pair of Berglund-Hübsch mirror elliptic curves $Z_{A,\langle J\rangle}=X_{A}$ and $Z_{A^{T},\langle J\rangle}=X_{A^{T}}$.
@@
Remark 4.3.
Conjecture 4.1 can be thought of as an arithmetic analogue of the Crepant Resolution Conjecture. In light of [Yasuda04], one may wonder whether a more general statement for Gorenstein orbifolds might hold. Since at present the scope of our methods is limited to the Calabi-Yau setting of Borisov’s construction, we leave this line of inquiry for future work.
@@
where $t^{\prime}$ is chosen to satisfy $t$ $^{\prime}\equiv t\text{ mod }p-1$ and $0<t^{\prime}\leq p-1$. Adapting the calculations of [Weil49] to the $p$-adic setting by employing the Techmuüller and Dwork characters introduced above, we have that the number of points of $Z_{A,G}(\mathbb{F}_{p})$ is equal to
@@
valid whenever $d$ divides p1p-1, and identifying indices by setting $\beta_{i}=A_{ii}-\gamma_{i}$, we obtain the equality of (4.2) and (4.4). ∎
@@
Table 1. Singular weighted diagonal K3 hypersurfaces. Multiplicity refers to the number of singular points of that type on $X_{A}$ . The last column lists the number of 2-2 curves appearing in the minimal resolution of the corresponding singularity
@@
Weighted diagonal K3 surfaces over finite fields are classified in [Goto]. Two of them, the Fermat quartic in $\mathbb{P}^{3}$ and the hypersurface $x_{1}^{2}+x_{2}^{6}+x_{3}^{6}+x_{4}^{6}=0$ in $\mathbb{P}(3,1,1,1)$ are smooth and Theorem 4.4 applies to them. The remaining $12$ cases with corresponding lists of singularities and the number of 2-2 curves added in the corresponding minimal resolution are listed in Table 1. In general, as observed in [Goto], some of the singularities in Table 1 are only defined over suitable field extensions. However, because in our case the coefficients in front of the monomials of $W_{A}(x)$ are always equal to $1$, these extensions are generated by elements $\gamma$ such that $1+\gamma^{m}=0$ with $m\in\{1,2\}$ depending on the singularity. On the other hand, inspection of Table 1 shows that $\det(A)$, and thus p1p-1, is always divisible by $4$. Hence our assumptions guarantee that $\mathbb{F}_{p}$ has primitive fourth roots of unity and thus $\mathbb{F}_{p}(\gamma)=\mathbb{F}_{p}$. Therefore, all the singularities in Table 1 actually occur, just as they do over $\mathbb{C}$, when $A$ satisfies the assumptions of Conjecture 4.1. Combining the calculations of [Weil49] with the (4.5) as in the proof of Theorem 4.4, we obtain
$X_{A}(\mathbb{F}_{p})=1+p+p^{2}+\sum_{\gamma}(-p)^{{\rm age}^{\vee}(\gamma)-1}\Gamma_{p}(\gamma A^{-1})\,$, (4.9)
@@
$Z_{A,\langle J\rangle}(\mathbb{F}_{p})=1+(\nu+1)p+p^{2}+\sum_{\gamma}(-p)^{{\rm age}^{\vee}(\gamma)-1}\Gamma_{p}(\gamma A^{-1})\,$, (4.10)
where $\nu$ is the total number of 2-2 curves appearing in $Z_{A,\langle J\rangle}$ viewed as the minimal resolution of $X_{A}$. Since the sum over $\gamma$ in (4.10) is equal to the contribution from terms labeled by $\lambda=0$ terms to (3.18), it remains to check that the contribution from the $\lambda\neq 0$ terms in (3.18) is exactly $1+(\nu+1)p+p^{2}$. It is easy to see that $y_{1}y_{2}y_{3}y_{4}$ contributes to 1, while $y_{1}^{A_{11}-1}$ where $\nu$ is the total number of 2-2 curves appearing in $Z_{A,\langle J\rangle}$ viewed as the minimal resolution of $X_{A}$. Since the sum over $\gamma$ in (4.10) is equal to the contribution from terms labeled by $\lambda=0$ terms to (3.18), it suffices to check that the contribution from the $\lambda\neq 0$ terms in (3.18) is exactly $1+(\nu+1)p+p^{2}$. It is easy to see that $y_{1}y_{2}y_{3}y_{4}$ contributes $1$ while $y_{1}^{A_{11}-1}y_{2}^{A_{22}-1}y_{3}^{A_{33}-1}y_{4}^{A_{44}-1}$ contributes $p^{2}$. By Remark 2.22, we know that the remaining $\lambda\neq 0$ terms all correspond to cohomology classes in $H^{1,1}(Z_{A,G}(\mathbb{C}))$. In particular, there must be precisely $\nu$ of them and they must be of one of two kinds: monomials of the form $y^{\lambda}$ with ${\rm age}(\lambda)=1$ and monomials of the form $x_{i}^{\gamma_{i}}x_{j}^{\gamma_{j}}y_{r}^{\lambda_{r}}y_{s}^{\lambda_{s}}\theta_{i}\theta_{j}$ with ${\rm age}(\lambda)=1={\rm age}^{\vee}(\gamma)$. Each of these $y^{\lambda}$ monomials contributes $p$ to (3.18). The constraint ${\rm age}^{\vee}(\gamma)=1$ forces each monomial $x_{i}^{\gamma_{i}}x_{j}^{\gamma_{j}}y_{r}^{\lambda_{r}}y_{s}^{\lambda_{s}}\theta_{i}\theta_{j}$ to be such that $\frac{\gamma_{i}}{A_{ii}}+\frac{\gamma_{j}}{A_{jj}}=1$, making the corresponding contribution to (3.18) equal to
$p\Gamma_{p}\left(\frac{\gamma_{i}}{A_{ii}}\right)\Gamma_{p}\left(\frac{\gamma_{j}}{A_{jj}}\right)=-p(-1)^{p-(p-1)\frac{\gamma_{i}}{A_{ii}}}\,$, (4.11)
where the reflection formula (3.22) was used to derive (4.11). Inspection of Table 1 shows that each $W_{A}(x)$ has at least two terms with even exponents. Taking again into account that $\det(A)$ divides p1p-1, we conclude that $\frac{p-1}{A_{ii}}$ is necessarily even and thus $p-(p-1)\frac{\gamma_{i}}{A_{ii}}$ is necessarily odd. Hence, monomials of the form $x_{i}^{\gamma_{i}}x_{j}^{\gamma_{j}}y_{r}^{\lambda_{r}}y_{s}^{\lambda_{s}}\theta_{i}\theta_{j}$ contribute by $p$ to (3.18), concluding the proof.
Remark 4.8.
The results of [Goto] require the assumption that $p$ is coprime to each of the weights of the ambient weighted projective space. In our case, this condition is automatically satisfied by our assumption that $\det(A)$ divides p1p-1. Indeed, the largest prime that appears as a factor of one of the weights in the list of weighted diagonal hypersurfaces in [Goto] is $7$, while the smallest value of $\det(A)$ in the same list is $256$.
@@
$\left(0,0\right)$ $y_{1}^{1}y_{2}^{1}y_{3}^{1}y_{4}^{1}$ 11
$\left(2,0\right)$ $x_{1}^{1}x_{2}^{1}x_{3}^{1}x_{4}^{1}\theta_{1}\theta_{2}\theta_{3}\theta_{4}$ $\Gamma_{p}\left(\frac{1}{2}\right)\Gamma_{p}\left(\frac{1}{3}\right)\Gamma_{p}\left(\frac{1}{10}\right)\Gamma_{p}\left(\frac{1}{15}\right)$